OpenStudy (anonymous):

(-3w^3 x^-6)^3 Write your answer using only positive exponents.

OpenStudy (anonymous):

\[x^{-a} = \frac{1}{x^a}\]\[(xy)^m = x^my^m\]

jimthompson5910 (jim_thompson5910):

how far did you get gabriel72alv?

OpenStudy (anonymous):

I couldnt figure out how to put it in the equation I needed to solve it? ):

jimthompson5910 (jim_thompson5910):

I'll do a similar example (just with different numbers) \[\Large \left(-2w^{5}x^{-8}\right)^{4}\] \[\Large \left(-2\right)^{4}\left(w^{5}\right)^{4}\left(x^{-8}\right)^{4}\] \[\Large \left((-2)^{1}\right)^{4}\left(w^{5}\right)^{4}\left(x^{-8}\right)^{4}\] \[\Large (-2)^{1*4}w^{5*4}x^{-8*4}\] \[\Large (-2)^{4}w^{20}x^{-32}\] \[\Large (-2)^{4}w^{20}\frac{1}{x^{32}}\] \[\Large 16w^{20}\frac{1}{x^{32}}\] \[\Large \frac{16w^{20}}{x^{32}}\]

OpenStudy (anonymous):

So would I get (-3w^3/x^6)^3 ?

jimthompson5910 (jim_thompson5910):

Keep going. So far, so good.

OpenStudy (anonymous):

Theres still another step? o:

jimthompson5910 (jim_thompson5910):

Yes there are more steps.

OpenStudy (anonymous):

How would you work it out to get to the next step though? I understand that.

jimthompson5910 (jim_thompson5910):

You would apply the outermost exponent to each term inside

jimthompson5910 (jim_thompson5910):

In other words, you would cube each piece.

OpenStudy (anonymous):

So basically would it be, (-3w^3/x^6) times (-3w^3/x^6) times (-3w^3/x^6) ?

jimthompson5910 (jim_thompson5910):

Yes pretty much

jimthompson5910 (jim_thompson5910):

So that explains why, for instance, in the denominator you will have x^6*x^6*x^6 = x^(6+6+6) = x^(18)

jimthompson5910 (jim_thompson5910):

A shortcut is to use this rule \[\Large (x^m)^n = x^{m*n}\]

OpenStudy (anonymous):

So would I get for my answer to that is -27w^9/x^18 ?

jimthompson5910 (jim_thompson5910):

You are 100% correct. Nice work.

OpenStudy (anonymous):

Nicee! Haa, So thats my answer? xD

jimthompson5910 (jim_thompson5910):

Yes it is.

OpenStudy (anonymous):

Awesome, Thank you so much jim_thompson5910 ! (:

jimthompson5910 (jim_thompson5910):

you're welcome